# Biomedical SAM 2: Segment Anything in Biomedical Images and Videos

### v1

@@ Abstract
Medical image segmentation and video object segmentation are essential for diagnosing and analyzing diseases by identifying and measuring biological structures. Recent advances in natural domain have been driven by foundation models like the Segment Anything Model 2 (SAM 2). To explore the performance of SAM 2 in biomedical applications, we designed two evaluation pipelines for single-frame image segmentation and multi-frame video segmentation with varied prompt designs, revealing SAM 2’s limitations in medical contexts. Consequently, we developed BioSAM 2, an enhanced foundation model optimized for biomedical data based on SAM 2. Our experiments show that BioSAM 2 not only surpasses the performance of existing state-of-the-art foundation models but also matches or even exceeds specialist models, demonstrating its efficacy and potential in the medical domain.
Figure 1: Overview Biomedical Tasks in BioSAM 2.
@@ 1 Introduction
Recognizing the need to extend these capabilities to more complex scenarios, SAM 2 [12] was developed to expand SAM’s functionality to include video inputs. This extension enables SAM 2 to process temporal sequences of images, making it suitable for tasks that require understanding spatial continuity over multiple frames. By handling both spatial and temporal dimensions, SAM 2 has demonstrated impressive zero-shot performance on various tasks involving natural image and video segmentation.
However, SAM 2’s potential on medical segmentation tasks has not yet been fully explored. We conducted a comprehensive evaluation to investigate its capability. Specifically, we assessed the performance of four SAM 2 variants (Hiera-T, Hiera-S, Hiera-B+, and Hiera-L) across 6 medical modalities and 9 objects of interest. We designed two evaluation pipelines for single-frame image segmentation and multi-frame video segmentation, respectively, incorporating diverse prompt designs. To further evaluate its performance, we compared it against various baseline models, including CNN-based, Transformer-based, and SSM-based models, using a range of metrics. Our findings indicate that SAM 2 cannot be directly utilized for medical image or video segmentation. The main reasons are the significant domain gap between natural and medical data and its inability to associate segmentation regions with meaningful semantic classes. In other words, SAM 2 cannot perform semantic segmentation in the medical domain, which limits its application in computer-aided diagnosis.
Building on these observations, we further introduce BioSAM 2, a refined foundation model that significantly enhances the segmentation performance of SAM 2 on biomedical images and videos. BioSAM 2 incorporates a memory mechanism and a stream processing architecture, which is the sam as SAM 2. This model can handle multi-frame segmentation tasks by retaining information from past predictions, allowing it to make accurate predictions on slices without explicit prompts. Experimental results demonstrate that BioSAM 2 consistently outperforms the state-of-the-art (SOTA) segmentation foundation model [6, 12], while on par with, or even surpassing the performance of specialist models [13, 14, 15, 16] trained on medical data from the same modality. These findings underscore the potential of BioSAM 2 as a new paradigm for versatile medical image and video segmentation.
Figure 2: Image segmentation results of tiny SAM 2 and large SAM 2 based on different segmentation prompts.
@@
  • We have developed two evaluation pipelines tailored for single-frame biomedical images and multi-frame biomedical videos within 6 medical modalities and 9 objects of interest. These pipelines comprehensively assess the performance of SAM 2 in biomedical applications.
  • To enhance the adaptability of SAM 2 in the biomedical domain, we introduce BioSAM 2, an optimized foundational model achieved by fine-tuning the original SAM 2. This refinement significantly boosts the segmentation performance of SAM 2. Remarkably, without any prompts, our automated segmentation consistently outperforms competitive SOTA foundation methods by a large margin.
  • The results of experiments show that BioSAM 2 matches or even exceeds the performance of specialized models trained on medical data of the same modality. These findings highlight the potential of BioSAM 2 as a novel paradigm for versatile medical applications.
By integrating BioSAM 2 with medical image segmentation tasks, we anticipate significant improvements in segmentation accuracy and annotation efficiency, ultimately contributing to better clinical outcomes and facilitating the adoption of AI in medical imaging. This research aims to push the boundaries of current medical image segmentation techniques and explore the full potential of advanced AI models like BioSAM 2 in handling the complexities of medical imaging data.
@@ 2 Related Work
The recently introduced SAM 2 [12] extended the backbone of SAM to 3D, enhancing its capability to “segment anything in videos”. Specifically, SAM 2 is equipped with a memory that stores information about objects and previous interactions, allowing it to generate and correct masklet predictions throughout the video.
@@ 3 Method
3.1 Preliminary Study of SAM 2
Segment Anything Model 2 (SAM 2) is a unified transformer-based model for both image and video segmentation. For each video frame, the segmentation prediction leverages the current prompt and previously observed memories. Videos are processed sequentially, with each frame handled individually by image encoder, while memory attention conditions current frame features on past frames and predictions. The mask decoder, which can optionally take input prompts, predicts the segmentation mask for that frame. Finally, a memory encoder transforms the predictions and image embeddings into a form usable for future frames, ensuring temporal consistency among multiple frames.
The vision transformer in the image encoder is pretrained using the hierarchical masked autoencoder model Hiera [42], enabling multiscale feature decoding. Memory attention conditions the current frame features on past frames’ features and predictions. Multiple transformer blocks are stacked, with the first block taking the current frame’s image encoding as input. Each block performs self-attention, followed by cross-attention to memories of frames and object pointers stored in a memory bank. In SAM 2, prompts are encoded with positional encoding and two learnable tokens specifying foreground and background. The mask decoder comprises bi-directional transformer blocks that update prompt and frame embeddings. The model predicts multiple masks per frame, and in cases of ambiguity without subsequent prompt clarification, it propagates only the mask with the highest predicted IoU. Additionally, an auxiliary prediction head determines the presence of the target object in the current frame. Finally, the memory encoder consolidates this process by downsampling the output mask using a convolutional module, which is then element-wise summed with the unconditioned frame embedding from the image encoder. This integrated data is retained in a memory bank, preserving essential information about past predictions for the target object throughout the video sequence.
Figure 3: The workflow of proposed BioSAM 2. We adapt SAM 2 for medical image and video segmentation by freezing the prompt encoder and only finetuning the image encoder and mask decoder.
3.2 Medical Applications of SAM 2
Building on the impressive zero-shot capabilities demonstrated by SAM 2 in natural images and videos, we explore its performance in medical applications. Specifically, we designed two pipelines, single-frame image segmentation and multi-frame video segmentation to assess SAM 2’s ability to handle zero-shot segmentation tasks within these medical contexts.
Single-frame Image Segmentation. We design single-frame image segmentation in a non-iterative manner in which all prompts are determined without feedback from any prior predictions.
@@ 3.2 Medical Applications of SAM 2
As highlighted in the SAM 2 documentation [12], employing a single point prompt can lead to segmentation ambiguity, as the model may associate the prompt with multiple valid masks without being able to discern the intended one. Although SAM 2 incorporates an ambiguity-resolving module that generates multiple masks and ranks them based on confidence scores, the use of multiple point prompts significantly mitigates this issue. Thus, we also assessed the performance of SAM 2 with additional randomly selected points from the initial mask candidates.
@@
Multi-frame Video Segmentation. In evaluating video segmentation, we employ a hybrid mode of offline and online evaluation. Specifically, we select the first n frames as interacted frames, on which click cues are added. Multiple click cues enable the model to more accurately determine the boundaries of objects, achieving higher segmentation accuracy. Additionally, multiple interacted frames effectively resolve the issue identified in SAM 2 where the model fails to track objects appearing in unmarked frames, making it more reasonable and effective for video scenarios. Overall, our evaluation traverses the video only once, subsequently yielding the final segmentation results.
@@
To summarize, when directly employing SAM 2 for medical image and video segmentation, the generated mask can be ambiguous and necessitates multiple prompts or iterations for prediction and correction. Despite its promising potential, SAM 2 has encountered challenges in delivering satisfactory segmentation results across various medical image/video segmentation tasks. Additionally, since the video training data for SAM 2 predominantly consists of high-resolution footage, it may become entirely ineffective when dealing with low-resolution medical videos. In light of these limitations, the objective of this study is to develop a robust segmentation foundation model capable of effectively addressing a wide range of segmentation targets.
3.3 BioSAM 2: Dedicated biomedical segmentation foundation model
@@
To tailor SAM 2 for medical video segmentation, selecting the appropriate network components for fine-tuning is crucial. The architecture of SAM 2 comprises several key elements: image encoder, prompt encoder, memory attention, mask decoder, and memory encoder. It is possible to fine-tune any combination of these components. For our adaptation, the prompt encoder, which processes the information from the given prompts, is retained from the pre-trained model and is therefore frozen to maintain its original functionality. Conversely, the image encoder and mask decoder are fine-tuned to enhance SAM 2’s suitability for the medical imaging domain. This adaptation strategy is depicted in Figure 3.
@@
For the image encoder, we opted for a tiny-sized configuration to strike a balance between computational cost and performance efficacy. We initiated training for SAM 2 from the official checkpoint, conducting separate sessions for image and video data. During training, we employed the AdamW optimizer [43] and implemented layer decay strategies [44] on the image encoder to enhance its learning efficiency. In terms of the mask decoder, its configuration is simplified to generate a single mask per frame, given that the provided prompt distinctly identifies the expected segmentation target.
@@
$\mathcal{L}=\mathcal{L}_{\text{dice}}+\mathcal{L}_{\text{BCE}}$ (4)
@@ 4.1.1 Datasets
To assess the performance and scalability of BioSAM 2, we utilize two medical image datasets, including Endoscopy dataset [45] and Microscopy dataset [46].
@@ 4.1.2 Experimental Setup
The setting of our experiments is the same as that in the corresponding official repository for each method to ensure a fair comparison. Specifically, the batch sizes for the Endoscopy and Microscopy datasets are 8. We adopt an unweighted combination of Dice loss and cross-entropy loss for all datasets and utilize the AdamW optimizer with an initial learning rate of 1e-4. The training duration for each dataset is set to 200 epochs. For the evaluation of SAM-based methods, we follow the implementation details of their official technique reports. We introduce different points as prompts and feed them into the model. All points are randomly chosen from the corresponding mask of the image.
@@
In our evaluation of BioSAM 2, we compare against two prominent CNN-based segmentation methods: nnUNet [13] and SegResNet [14]. Additionally, we include a comparison with UNETR [15] and SwinUNETR [16], two Transformer-based networks that have gained popularity in medical image segmentation tasks. U-Mamba [22], a recent method based on the Mamba model, is also included in our comparison to provide a comprehensive overview of its performance. For each model, we implement their recommended optimizers to ensure consistency in training conditions. To maintain fairness across all comparisons, we finetune all these models on each dataset and apply the default image preprocessing in nnUNet [13]. We also assess the performance of SAM and SAM 2 by directly allowing them to infer the corresponding mask for the image. To ensure a thorough evaluation, we test two sizes of SAM 2 using three kinds of prompt.
@@
For evaluation metrics, we employ the Dice Similarity Coefficient (DSC) and Normalized Surface Distance (NSD) to assess performance in instrument segmentation in endoscopy images. For the cell segmentation task, we utilize the F1 score and NSD to evaluate method performance.
@@
Table 1: Results summary of image segmentation tasks: instruments segmentation in endoscopy images, and cell segmentation in microscopy images.
@@
Methods Instruments in Endoscopy Cells in Microscopy DSC NSD F1 NSD nnU-Net 0.6264$\pm$0.3024 0.6412$\pm$0.3074 0.5383$\pm$0.2657 0.8332$\pm$0.1611 SegResNet 0.5820$\pm$0.3268 0.5968$\pm$0.3303 0.5411$\pm$0.2633 0.7944$\pm$0.2356 UNETR 0.5017$\pm$0.3201 0.5168$\pm$0.3235 0.4357$\pm$0.2572 0.8201$\pm$0.2263 SwinUNETR 0.5811$\pm$0.3176 0.5973$\pm$0.3209 0.3880$\pm$0.2664 0.7981$\pm$0.1920 U-Mamba 0.6303$\pm$0.3067 0.6451$\pm$0.3104 0.5607$\pm$0.2784 0.8288$\pm$0.1706 SAM 0.1583$\pm$0.3300 0.1600$\pm$0.3313 0.3249$\pm$0.2285 0.3696$\pm$0.2544 SAM 2 (Hiera-T, 1-click) 0.4115$\pm$0.4092 0.4227$\pm$0.4189 0.0654$\pm$0.1220 0.0720$\pm$0.1372 SAM 2 (Hiera-T, 3-click) 0.5215$\pm$0.3802 0.5349$\pm$0.3864 0.3436$\pm$0.2400 0.3911$\pm$0.2640 SAM 2 (Hiera-T, 5-click) 0.5382$\pm$0.3568 0.5520$\pm$0.3616 0.3566$\pm$0.2496 0.4070$\pm$0.2761 SAM 2 (Hiera-L, 1-click) 0.4416$\pm$0.4217 0.4523$\pm$0.4312 0.0799$\pm$0.1540 0.0877$\pm$0.1722 SAM 2 (Hiera-L, 3-click) 0.5354$\pm$0.3750 0.5497$\pm$0.3813 0.3217$\pm$0.2481 0.3720$\pm$0.2750 SAM 2 (Hiera-L, 5-click) 0.5479$\pm$0.3629 0.5623$\pm$0.3681 0.3352$\pm$0.2598 0.3876$\pm$0.2849 BioSAM 2 (Hiera-T) 0.6251$\pm$0.2897 0.6427$\pm$0.3095 0.5792$\pm$0.2666 0.7436$\pm$0.2104

### v2

# Biomedical SAM-2: Segment Anything in Biomedical Images and Videos

@@ Abstract
Medical image segmentation and video object segmentation are essential for diagnosing and analyzing diseases by identifying and measuring biological structures. Recent advances in natural domain have been driven by foundation models like the Segment Anything Model 2 (SAM-2). To explore the performance of SAM-2 in biomedical applications, we designed three evaluation pipelines for single-frame 2D image segmentation, multi-frame 3D image segmentation and multi-frame video segmentation with varied prompt designs, revealing SAM-2’s limitations in medical contexts. Consequently, we developed BioSAM-2, an enhanced foundation model optimized for biomedical data based on SAM-2. Our experiments show that BioSAM-2 not only surpasses the performance of existing state-of-the-art foundation models but also matches or even exceeds specialist models, demonstrating its efficacy and potential in the medical domain.
Figure 1: Overview of biomedical tasks in BioSAM-2.
@@ 1 Introduction
Recognizing the need to extend these capabilities to more complex scenarios, SAM-2 [12] was developed to expand SAM’s functionality to include video inputs. This extension enables SAM-2 to process temporal sequences of images, making it suitable for tasks that require understanding spatial continuity over multiple frames. By handling both spatial and temporal dimensions, SAM-2 has demonstrated impressive zero-shot performance on various tasks involving natural image and video segmentation.
However, SAM-2’s potential on medical segmentation tasks has not yet been fully explored. We conducted a comprehensive evaluation to investigate its capability. Specifically, we assessed the performance of four SAM-2 variants (Hiera-T, Hiera-S, Hiera-B+, and Hiera-L) across 8 medical modalities and 22 objects of interest. We designed three evaluation pipelines for single-frame 2D image segmentation, multi-frame 3D image segmentation and multi-frame video segmentation, respectively, incorporating diverse prompt designs. To further evaluate its performance, we compared it against various baseline models, including CNN-based, Transformer-based, and SSM-based models, using a range of metrics. Our findings indicate that SAM-2 cannot be directly utilized for medical image or video segmentation. The main reasons are the significant domain gap between natural and medical data and its inability to associate segmentation regions with meaningful semantic classes. In other words, SAM-2 cannot perform semantic segmentation in the medical domain, which limits its application in computer-aided diagnosis.
Building on these observations, we further introduce BioSAM-2, a refined foundation model that significantly enhances the segmentation performance of SAM-2 on biomedical images and videos. BioSAM-2 incorporates a memory mechanism and a stream processing architecture, which is the sam as SAM-2. This model can handle multi-frame segmentation tasks by retaining information from past predictions, allowing it to make accurate predictions on slices without explicit prompts. Experimental results demonstrate that BioSAM-2 consistently outperforms the state-of-the-art (SOTA) segmentation foundation model [6, 12], while on par with, or even surpassing the performance of specialist models [13, 14, 15, 16] trained on medical data from the same modality. These findings underscore the potential of BioSAM-2 as a new paradigm for versatile medical image and video segmentation.
Figure 2: Image segmentation results of tiny SAM-2 and large SAM-2 based on different segmentation prompts.
@@
  • We have developed three evaluation pipelines tailored for single-frame biomedical images, multi-frame biomedical images and multi-frame biomedical videos within 8 medical modalities and 22 objects of interest. These pipelines comprehensively assess the performance of SAM-2 in biomedical applications.
  • To enhance the adaptability of SAM-2 in the biomedical domain, we introduce BioSAM-2, an optimized foundational model achieved by fine-tuning the original SAM-2. This refinement significantly boosts the segmentation performance of SAM-2. Remarkably, without any prompts, our automated segmentation consistently outperforms competitive SOTA foundation methods by a large margin.
  • The results of experiments show that BioSAM-2 matches or even exceeds the performance of specialized models trained on medical data of the same modality. These findings highlight the potential of BioSAM-2 as a novel paradigm for versatile medical applications.
By integrating BioSAM-2 with medical image segmentation tasks, we anticipate significant improvements in segmentation accuracy and annotation efficiency, ultimately contributing to better clinical outcomes and facilitating the adoption of AI in medical imaging. This research aims to push the boundaries of current medical image segmentation techniques and explore the full potential of advanced AI models like BioSAM-2 in handling the complexities of medical imaging data.
@@ 2 Related Work
The recently introduced SAM-2 [12] extended the backbone of SAM to 3D, enhancing its capability to “segment anything in videos”. Specifically, SAM-2 is equipped with a memory that stores information about objects and previous interactions, allowing it to generate and correct masklet predictions throughout the video.
@@ 3 Method
3.1 Preliminary Study of SAM-2
Segment Anything Model 2 (SAM-2) is a unified transformer-based model for both image and video segmentation. For each video frame, the segmentation prediction leverages the current prompt and previously observed memories. Videos are processed sequentially, with each frame handled individually by image encoder, while memory attention conditions current frame features on past frames and predictions. The mask decoder, which can optionally take input prompts, predicts the segmentation mask for that frame. Finally, a memory encoder transforms the predictions and image embeddings into a form usable for future frames, ensuring temporal consistency among multiple frames.
The vision transformer in the image encoder is pretrained using the hierarchical masked autoencoder model Hiera [42], enabling multiscale feature decoding. Memory attention conditions the current frame features on past frames’ features and predictions. Multiple transformer blocks are stacked, with the first block taking the current frame’s image encoding as input. Each block performs self-attention, followed by cross-attention to memories of frames and object pointers stored in a memory bank. In SAM-2, prompts are encoded with positional encoding and two learnable tokens specifying foreground and background. The mask decoder comprises bi-directional transformer blocks that update prompt and frame embeddings. The model predicts multiple masks per frame, and in cases of ambiguity without subsequent prompt clarification, it propagates only the mask with the highest predicted IoU. Additionally, an auxiliary prediction head determines the presence of the target object in the current frame. Finally, the memory encoder consolidates this process by downsampling the output mask using a convolutional module, which is then element-wise summed with the unconditioned frame embedding from the image encoder. This integrated data is retained in a memory bank, preserving essential information about past predictions for the target object throughout the video sequence.
Figure 3: The workflow of proposed BioSAM-2. We adapt SAM-2 for medical image and video segmentation by freezing the prompt encoder and only finetuning the image encoder and mask decoder.
3.2 Medical Applications of SAM-2
Building on the impressive zero-shot capabilities demonstrated by SAM-2 in natural images and videos, we explore its performance in medical applications. Specifically, we designed three pipelines, single-frame image segmentation, multi-frame iamge segmentation and multi-frame video segmentation to assess SAM-2’s ability to handle zero-shot segmentation tasks within these medical contexts.
Single-frame 2D Image Segmentation. We design single-frame image segmentation in a non-iterative manner in which all prompts are determined without feedback from any prior predictions.
@@ 3.2 Medical Applications of SAM-2
As highlighted in the SAM-2 documentation [12], employing a single point prompt can lead to segmentation ambiguity, as the model may associate the prompt with multiple valid masks without being able to discern the intended one. Although SAM-2 incorporates an ambiguity-resolving module that generates multiple masks and ranks them based on confidence scores, the use of multiple point prompts significantly mitigates this issue. Thus, we also assessed the performance of SAM-2 with additional randomly selected points from the initial mask candidates.
@@
Multi-frame 3D Image Segmentation. For multi-frame 3D image segmentation, we initially set the bounding box prompts on the middle slice of the image volume, following [43, 44]. The middle slice is chosen as the starting point in 3D image segmentation due to its tendency to contain the largest object and the greatest number of semantic classes among all slices in the conventional axial view. This strategic choice enhances the accuracy and reliability of subsequent segmentation. Once the middle frame is annotated, SAM-2 facilitates the propagation of annotations to the surrounding slices that do not have prompts. This is achieved through the video segmentation feature integrated into SAM-2. The process involves first propagating the annotations backward from the middle slice to the first slice, followed by a forward propagation from the middle slice to the last slice in the volume. This bidirectional propagation ensures consistent segmentation across the entire image volume.
@@
Multi-frame Video Segmentation. In evaluating video segmentation, we employ a hybrid mode of offline and online evaluation. Specifically, we select the first n frames as interacted frames, on which click cues are added. Multiple click cues enable the model to more accurately determine the boundaries of objects, achieving higher segmentation accuracy. Additionally, multiple interacted frames effectively resolve the issue identified in SAM-2 where the model fails to track objects appearing in unmarked frames, making it more reasonable and effective for video scenarios. Overall, our evaluation traverses the video only once, subsequently yielding the final segmentation results.
@@
To summarize, when directly employing SAM-2 for medical image and video segmentation, the generated mask can be ambiguous and necessitates multiple prompts or iterations for prediction and correction. Despite its promising potential, SAM-2 has encountered challenges in delivering satisfactory segmentation results across various medical image/video segmentation tasks. Additionally, since the video training data for SAM-2 predominantly consists of high-resolution footage, it may become entirely ineffective when dealing with low-resolution medical videos. In light of these limitations, the objective of this study is to develop a robust segmentation foundation model capable of effectively addressing a wide range of biomedical segmentation targets.
3.3 BioSAM-2: Dedicated biomedical segmentation foundation model
@@
To tailor SAM-2 for medical video segmentation, selecting the appropriate network components for fine-tuning is crucial. The architecture of SAM-2 comprises several key elements: image encoder, prompt encoder, memory attention, mask decoder, and memory encoder. It is possible to fine-tune any combination of these components. For our adaptation, the prompt encoder, which processes the information from the given prompts, is retained from the pre-trained model and is therefore frozen to maintain its original functionality. Conversely, the image encoder and mask decoder are fine-tuned to enhance SAM-2’s suitability for the medical imaging domain. This adaptation strategy is depicted in Figure 3.
@@
For the image encoder, we opted for a tiny-sized configuration to strike a balance between computational cost and performance efficacy. We initiated training for SAM-2 from the official checkpoint, conducting separate sessions for image and video data. During training, we employed the AdamW optimizer [45] and implemented layer decay strategies [46] on the image encoder to enhance its learning efficiency. In terms of the mask decoder, its configuration is simplified to generate a single mask per frame, given that the provided prompt distinctly identifies the expected segmentation target.
@@
$\mathcal{L}=\alpha\mathcal{L}_{\text{dice}}+\beta\mathcal{L}_{\text{BCE}}$ (4)
@@ 4.1.1 Datasets
To assess the performance and scalability of BioSAM-2, we utilize four medical image datasets, including Abdomen CT dataset [47], Abdomen MR dataset [48], Endoscopy dataset [49] and Microscopy dataset [50].
Abdomen 3D images.  For evaluating BioSAM-2, two datasets are implemented. The first is Abdomen CT [47] from the MICCAI 2022 FLARE challenge, which includes segmentation of 13 abdominal organs from 50 CT scans in both the training and testing sets. The organs include the liver, spleen, pancreas, kidneys, stomach, gallbladder, esophagus, aorta, inferior vena cava, adrenal glands, and duodenum. The second dataset, Abdomen MR [48] from the MICCAI 2022 AMOS Challenge, involves the same 13 organs with 60 MRI scans for training and 50 for testing.
@@ 4.1.2 Experimental Setup
The setting of our experiments is the same as that in the corresponding official repository for each method to ensure a fair comparison. We adopt an unweighted combination of Dice loss and cross-entropy loss for all datasets and utilize the AdamW optimizer with an initial learning rate of 1e-4. The training duration for each dataset is set to 200 epochs, with batch size as 8. For the evaluation of SAM-based methods, we follow the implementation details of their official technique reports. We introduce different prompt designs and feed them into the model. All prompts are randomly chosen from the corresponding mask of the image.
@@
In our evaluation of BioSAM-2, we compare against two prominent CNN-based segmentation methods: nnUNet [13] and SegResNet [14]. Additionally, we include a comparison with UNETR [15] and SwinUNETR [16], two Transformer-based networks that have gained popularity in medical image segmentation tasks. U-Mamba [22], a recent method based on the Mamba model, is also included in our comparison to provide a comprehensive overview of its performance. For each model, we implement their recommended optimizers to ensure consistency in training conditions. To maintain fairness across all comparisons, we finetune all these models on each dataset and apply the default image preprocessing in nnUNet [13]. We also assess the performance of SAM and SAM-2 by directly allowing them to infer the corresponding mask for the image. To ensure a thorough evaluation, we test two sizes of SAM-2 using three kinds of prompt.
@@
For evaluation metrics, we employ the Dice Similarity Coefficient (DSC) and Normalized Surface Distance (NSD) to assess performance in abdominal multi-organ segmentation for CT and MR scans, as well as instrument segmentation in endoscopy images. For the cell segmentation task, we utilize the F1 score and NSD to evaluate method performance.
@@
Table 1: Results summary of 2D image segmentation tasks: instruments segmentation in endoscopy images, and cell segmentation in microscopy images.
@@
Methods Instruments in Endoscopy Cells in Microscopy DSC NSD F1 NSD nnU-Net 0.6264$\pm$0.3024 0.6412$\pm$0.3074 0.5383$\pm$0.2657 0.8332$\pm$0.1611 SegResNet 0.5820$\pm$0.3268 0.5968$\pm$0.3303 0.5411$\pm$0.2633 0.7944$\pm$0.2356 UNETR 0.5017$\pm$0.3201 0.5168$\pm$0.3235 0.4357$\pm$0.2572 0.8201$\pm$0.2263 SwinUNETR 0.5811$\pm$0.3176 0.5973$\pm$0.3209 0.3880$\pm$0.2664 0.7981$\pm$0.1920 U-Mamba 0.6303$\pm$0.3067 0.6451$\pm$0.3104 0.5607$\pm$0.2784 0.8288$\pm$0.1706 SAM 0.1583$\pm$0.3300 0.1600$\pm$0.3313 0.3249$\pm$0.2285 0.3696$\pm$0.2544 SAM-2 (Hiera-T, 1-click) 0.4115$\pm$0.4092 0.4227$\pm$0.4189 0.0654$\pm$0.1220 0.0720$\pm$0.1372 SAM-2 (Hiera-T, 3-click) 0.5215$\pm$0.3802 0.5349$\pm$0.3864 0.3436$\pm$0.2400 0.3911$\pm$0.2640 SAM-2 (Hiera-T, 5-click) 0.5382$\pm$0.3568 0.5520$\pm$0.3616 0.3566$\pm$0.2496 0.4070$\pm$0.2761 SAM-2 (Hiera-L, 1-click) 0.4416$\pm$0.4217 0.4523$\pm$0.4312 0.0799$\pm$0.1540 0.0877$\pm$0.1722 SAM-2 (Hiera-L, 3-click) 0.5354$\pm$0.3750 0.5497$\pm$0.3813 0.3217$\pm$0.2481 0.3720$\pm$0.2750 SAM-2 (Hiera-L, 5-click) 0.5479$\pm$0.3629 0.5623$\pm$0.3681 0.3352$\pm$0.2598 0.3876$\pm$0.2849 BioSAM-2 (Hiera-T) 0.6251$\pm$0.2897 0.6427$\pm$0.3095 0.5792$\pm$0.2666 0.7436$\pm$0.2104